OpenStudy (anonymous):

The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x? x = 6 x = 4 x = -3 x = -2

OpenStudy (anonymous):

put the points into the slope formula y2-y1/x2-x1

OpenStudy (anonymous):

oh no sorry the midpoint formula :)

OpenStudy (anonymous):

i kinda just want the anwer dont have time to work it out

OpenStudy (austinl):

\[Midpoint = (\frac{x_1+x_2}{2} , \frac{y_1+y_2}{2})\]

OpenStudy (anonymous):

midpoint formula is (x1+x2/2,y1+y2/2)

OpenStudy (anonymous):

dont be lazy come on :)

OpenStudy (anonymous):

im doing alot of work so i can finish my class today and get my deploma on friday

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

@austinL if u then divide that by 2 u get 1, which is not the point

OpenStudy (austinl):

You are correct @groges. Props. I forgot the divide by two.

OpenStudy (anonymous):

the x value is given to be 2, which means in the formula it is x-2/2 =2 x-2=4 x=6

OpenStudy (anonymous):

thank you guys for the help